OpenStudy (nikkibxox):

Multiple Questions Algebra 2 a Which of the following is equal to (2x/5 -9)+9 Explain. A) 2x/5 B) 5x/2 C) 2x-45/ 5 D) (2x-9) +45 Solve and Explain; 2|x+2|-1=5 Solve and Explain; |-4b-8|+|-1-b^2|+ 2b^3; B=-2 Solve and Check for extanuos solutions then explain; 9|9=8x|=2x+3 What is f(-5) if F(x)=|2x-1|+10; Solve and Explain Write the equation -5x-12y=11 in slope-intercept form. What is the slope and Y-intercept. Explain.

OpenStudy (nikkibxox):

I will give medals! Please Help!

OpenStudy (texaschic101):

1. it would be 2x/5 because the 9's cancel each other out. ================================= 2. 2 | x + 2 | - 1 = 5 --- add 1 to both sides 2 | x + 2 | = 6 --- divide both sides by 2 | x + 2 | = 3.....we basically wanted to get the absolute value and everything in it by itself and everything else on the other side of the equal sign. Now we can set up 2 equations. x + 2 = 3 -(x + 2) = 3 x = 3 - 2 -x - 2 = 3 x = 1 -x = 5 x = -5 so...x = 1 and x = -5 ================================= 3. | -4b - 8 | + | -1 - b^2 | + 2b^3.....when b = -2 first we sub in our -2 for b | -4(-2) - 8 | + | -1 - (-2^2) | + 2(-2^3) = | 8 - 8 | + | -1 - (-2^2) | + 2(-2^3) = 0 + |- 1 - 4 | + 2(-8) = 0 + | -5 | - 16 =.....absolute values are positive even when they say negative 0 + 5 - 16 = - 11 You might need a second opinion on this one ================================= 4. ??...sorry ================================= 5. f(x) = | 2x - 1 | + 10.....when x = -5 sub in -5 for x f(-5) = | 2(-5) - 1 | + 10 f(-5) = | -10 - 1| + 10 f(-5) = | -11 | + 10....absolute values are positive f(-5) = 11 + 10 f(-5) = 21 ================================== 6. -5x - 12y = 11 .... add 5x to both sides -12y = 5x + 11 --- divide both sides by -12 y = (-5/12)x - 11/12 --- slope intercept form Now, in y = mx + b form, the slope is in the m position and the y intercept is in the b position. So your slope is -5/12 and your y int is -11/12

OpenStudy (nikkibxox):

Oh Mer Gerd. Your the best! Thank you!